La presentazione è in caricamento. Aspetta per favore

La presentazione è in caricamento. Aspetta per favore

RPC Barrel Status del Commissioning P. Vitulo Bari/Frascati/Napoli/Pavia Sinossi della presentazione.

Presentazioni simili


Presentazione sul tema: "RPC Barrel Status del Commissioning P. Vitulo Bari/Frascati/Napoli/Pavia Sinossi della presentazione."— Transcript della presentazione:

1 RPC Barrel Status del Commissioning P. Vitulo Bari/Frascati/Napoli/Pavia Sinossi della presentazione

2 Overview Stato dellinstallazione: cabling, elettronica Stato del commissioning: Tower test TTU test Gas P5 Closed Gas Gain Monitoring Sommario Commissioning Addenda: Report on gas leak (from A. Colaleo) Report on Power System (P. Paolucci) Report on Gas Gain Monitor (S. Bianco) Sinossi della presentazione

3 PV-8 maggio 2008 UXC Hardware BARREL Status WheelDetect-tower cables (HV-LV-DCS cables) Link Boxes Link Boards LV easy crates LV modules RBCTower cables MAO W+2 ok W+1 ok W0 ok W-1 ok W-2 ok 50% ok Overall Status: Missing half LV modules for W-2 LV modules expected in April/May at CERN Status installazione:cabling, elettronica

4 PV-8 maggio 2008 WheelHV cables –patch panel HV Easy Crates HV distrib. EASY – majiste SY1527 local cable HV multicond. Cable test Trigger cables from UXC-USC W+2 ok W+1 ok W0 ok W-1 ok W-2 ok USC55 hardware BARREL Status Overall status: HV modules not all in place Action:We move from one easy crates to the other in order to test the chambers. All missing modules will be at CERN in April/May First delivery phase: HV modules were affected by offset instability Second delivery phase (middle of March): HV modules were affected by negative offsets. Action: we solved the problem by software. Status installazione:cabling, elettronica

5 Steps towards full system commissioning: Commissioning steps Commissioning steps 1.Local commissioning from chamber to tower : Tower tests with random trigger 2.Commissioning with cosmics: test with RBC-TTU, efficiency studies (FEB threshold and HV working point ) 2.Commissioning of the full detector/DAQ/Trigger chain Global Runs/Minidaq Status Commissioning

6 PV-8 maggio 2008 Tower Test done for 9/10 towers so far W-2 Tower Test finished end of April 2008 Several inversions of signal cables, which were swapped either in LB or on CH side: solved 1-2 partition/ wheel not acknowledged (no threshold control, feb set at default threshold=200 mV) noisy FEB (~ % / tower) dead strips (~ % / tower) All problems reported in a web page. Complete map of all hardware problems: shift operator must check it before operate the chamber. Tower test summary Status Commissioning

7 PV-8 maggio 2008 Installation of RBC boards on the detector towers and cabling: done Installation of TTU boards in USC55: 1 out of 3 (can cover about 2 wheels) The installation will be completed as soon as full production of Trigger Boards (Warsaw) will be delivered Validation of implemented algorithm with local readout and minidaq. Test started on W+1 in March 1. At sector level 2.Tower level (Only with sect for the moment) 3.Wheel level (need full TBs system installed) Study the timing calibration for cosmic muons Xcheck with DT trigger Coupling with Global Trigger Provide Trigger for any cosmic run, May Global Run and CMS startup TTU Trigger status TTU status

8 PV-8 maggio 2008 WHV test UXC-USC LV, DCS test Tower testcosmic test with RBC-TTU (sect ) W+2ok May W+1ok March-April W0ok April-May W-1ok June W-2okFar side by middle of April June Summary of Commissioning status and plan Summary of Commissioning status and plan Status commissioning e pianificazione

9 PV-8 maggio 2008 Status del commissioning del P5 Il sistema finale funziona ed è stabile sulla W0. Dal 20 aprile aggiunta la W2 (in entrabi i casi e solo per ragioni di costi, ci stiamo limitando alla torre near: 6 settori). Le camere della W0 sono ON (sul sistema finale) dal 10 aprile. Previsto un sistema di interlock collegato al sistema di HV in caso di stop del sistema di gas. La qualità della miscela è costantemente monitorata con: GC, misuratore di O2 e H2O la concentrazione di aria nella miscela che torna dalle camere è a livelli piuttosto bassi L'umidificatore è in funzione e stiamo monitorando l'umidità del gas sia dopo l'umidificatore, sia a livello del distributore sulla W0 (2 canali hanno sensori di umidità sulla linea di supply e su quella di return). L'umidità relativa della miscela è 40%. La temperatura è di circa 20 C. Il flusso per canale di almeno 15 l/h. Entrambi i filtri sono in uso: la miscela che torna dalle camere passa attraverso i filtri e va in aria per ora (open loop). Il gas viene analizzato regolarmente prima e dopo i filtri. Se non ci saranno maggiori intoppi, inizieremo il ricircolo verso la metà di maggio (dopo circa un mese di run in open mode). Partiremo con una frazione di miscela nuova dell'ordine del 40-50%. Status of Gas System

10 PV-8 maggio 2008 Status of Gas System ISR (CERN Gas Group (Roberto Guida) + Gruppo RPC) Le correnti stanno aumentando (i filtri non sono mai stati rigenerati dallo scorso luglio). L'aumento è stato molto graduale, non repentino come la scorsa volta La concentrazione di aria nel loop è sostanzialmente uguale a quella della miscela nuova. Con il GC non si vedono particolari extra-componenti. Sono stati predisposti diversi gorgogliatori per il campionamento del gas in soluzione. La stazione per la misura dei fluoruri è pronta e funzionante. finite le misure in argon, ripartiremo con filtri nuovi (in particolare sostituiamo il setaccio molecolare 5A con il 3A). nel nuovo ciclo, saranno collegate anche le 10 Single gap nuove ed anche tre delle camerette del "gas gain monitoring". GAS system ISR

11 PV-8 maggio 2008 I rivelatori del Gas Gain Monitoring sono in operazione allISR, in corso lintegrazione DCS/PVSS/xDAQ Analisi preliminari della miscela durante i test di ricircolo hanno mostrato la presenza in tracce di elementi metallici, probabilmenti rilasciati dai filtri durante o dopo la saturazione Differenti procedure per lo studio della cinetica chimica dei filtri. Sebbene i filtri debbano essere caratterizzati chimicamente il setup utilizzato è capace di operare correttamente nel detector. GAS Gain monitor e ISR Status of Gas System

12 PV-8 maggio 2008 Steps for each wheel: Flow cell equalization 100% mixture in fill mode operation (6 sectors) with current and noise monitoring 50% mixture in closed mode (6 sectors) for 1 month with current and noise monitoring GAS system commissioning plan Goal : 1/2 Barrel (6 sectors/wheel) in recirculation mode (50% fresh gas) by end of May. All barrel wheels in recirculation mode (50% fresh gas) by end of July. Status of Gas System

13 PV-8 maggio 2008 All hardware (HV/LV) in place by end of April Trigger system ready for all wheels by end of May (now 50% installed) Tower test done on 9/10 towers. All hardware chain fully commissioned by May. 3 main hardware problems (chamber LHC start up: 2 chambers with big gas leak ( W0/RB3/S8, W+2/RB4/S8 ) 1 chamber with LV problem ( W-1/RB4/S6 ) Action: Chambers to be extracted and checked at first shutdown. 4 chambers with gas leaks but OPERABLE ( W0/RB1/S1, W0/RB3/S2, W-2/RB1/S7, W+1 RB4 ) Gas leak tests done three times on all chambers after installation. GAS system commissioning will require a lot of effort since also chamber performance must be checked in parallel. No enough experts available. Goal is to have full system fully commissioned in re-circulation mode (high % fresh gas) by July. Conclusion and status at startup Sommario Commissioning

14 PV-8 maggio settimane costruzione camere al CERN (x 3 tecnici ) appena possibile 2 settimana installazione camere da estrarre (x 3 tecnici) alla riapertura 2 settimane riparazione eventuali (cavi, connettori,moduli): x 3 tecnici shutdown 2 settimane riparazioni eventuali (accesso possibile): x 2 tecnici presa dati 2 settimane smantellamento ISR e Bd904 e spostamento a PT5(smantellamento ISR già stadio avanzato): x 3 tecnici entro dicembre Presenza costante di un tecnico equivalente multi-task Manpower - Tecnici Manpower per completamento e mantenimento Gas Gain Monitoring (assistenza prelievi all'ISR, disinstallazione dall'ISR, trasporto al SG5, installazione al SG5, checkout e assistenza commissioning camerette GGM ) 3 mu 5 mu (probabili overlap ma non tanti) mu (installazione) + 3 mu (gas Mon) TOT 9.5 mu

15 Manpower per SHIFT Manpower - Fisici RPC BARREL SYSTEM AREA DATA TAKING SHUT-DOWN MANAGERS OPERATORS EXPERTS TOTAL MANAGERS OPERATORS EXPERTS TOTAL TOTAL OPERATION 3,04,03,010,01,00,05,06,0 TOTAL MAINTENANCE 0,0 3,0 0,0 TOTAL SYSTEM AREA 3,04,06,013,01,00,05,06,0 TOT : 13 (mu /mesi RUN) + (6 mu/mesi Shutdown) 11 (mu /mesi RUN) + (6 mu/mesi Shutdown) a regime


Scaricare ppt "RPC Barrel Status del Commissioning P. Vitulo Bari/Frascati/Napoli/Pavia Sinossi della presentazione."

Presentazioni simili


Annunci Google